OpenStudy (anonymous):

x/8 - 5 = -1 Check my answer?

OpenStudy (anonymous):

I got x = 32

OpenStudy (anonymous):

x/8-5=-1 x/8=-1+5 x=8 * 4 so x = 32 @greenlegodude57 :)